Consider a gas with molar mass M.If the sound at frequency f is introduced to a tube of this gas at temperature T then an internal acoustic standing wave is set up with nodes separated by L.The adiabatic constant (gamma=Cp/ Cv is

Ans:(4Mf
1
Expert's answer
2021-08-25T12:25:30-0400

"v=2Lf,"

also "v=\sqrt{\frac{\gamma RT}M}," "\implies"

"4L^2f^2=\frac{\gamma RT}M,"

"\gamma=\frac{4Mf^2L^2}{RT}."